A ray of light is incident on a plane mirror along a vector  i cap + j cap +k cap . The normal on incidence point is along icap+ jcap . The unit vector along the reflected ray is?                                                   

Grade:11

1 Answers

Vikas TU
7 years ago
First write down the angle b/w the incident and normal and equate it to the angle of normal and reflected light.
Assume reflected light be ai + bj + ck.
its magnitude would be 1.
Therefore,
(a+b)/root(2) = (1+1)/root2*root3 = 2/root(6)
thus,
a+b = 2/root(3).
Now do same thing with plane line eqn. that would be perpendicular to normal and that is: icap – jcap.
but keep the angle this time 90 – thetha.